# Music Grounding by Short Video

### v2

@@ Abstract
Adding proper background music helps complete a short video to be shared. Previous research tackles the task by video-to-music retrieval (V2MR), which aims to find the most suitable music track from a collection to match the content of a given query video. In practice, however, music tracks are typically much longer than the query video, necessitating (manual) trimming of the retrieved music to a shorter segment that matches the video duration. In order to bridge the gap between the practical need for music moment localization and V2MR, we propose a new task termed Music Grounding by Short Video (MGSV). To tackle the new task, we introduce a new benchmark, MGSV-EC, which comprises a diverse set of 53k short videos associated with 35k different music moments from 4k unique music tracks. Furthermore, we develop a new baseline method, MaDe, which performs both video-to-music matching and music moment detection within a unifed end-to-end deep network. Extensive experiments on MGSV-EC not only highlight the challenging nature of MGSV but also sets MaDe as a strong baseline. Data and code will be released.
Figure 1: Music grounding by short video (MGSV), aiming to localize within a music-track collection a music moment that best serves as background music for the query video. The text in the callout boxes to the right of each music moment is for illustrative purposes only.
@@ 1 Introduction
To address the new task, one might consider applying V2MR at a fine-grained segment level, with music tracks pre-trimmed into shorter segments. This approach is problematic due to the varying durations of videos. While MGSV is new, we observe in a broader context its conceptual resemblance to Video Grounding (VG) [16, 14, 30], which aims to localize a moment in a specific video w.r.t. a given textual query. In particular, the query and target modalities in MGSV (VG) are video (text) and music (video), respectively. Thus, a VG method could, in principle, be repurposed for MGSV in a single-music mode, assuming that the relevant music track for a query video is provided. When dealing with a collection of music tracks, V2MR is required to identify the relevant track. While combining existing VG and V2MR methods provides a good starting point, these methods were not originally designed for MGSV. By incorporating task-specific optimizations, we develop a stronger baseline that performs video-to-music Matching and music moment Detection (MaDe) in a single network.
While few public datasets (mainly HIMV [11]) exist for V2MR, each video in these datasets is paired with an equal-length music segment, yet the source tracks from which these segments were derived are absent. Therefore, they cannot be repurposed to MGSV. We construct a new dataset, MGSV-EC, sourced from an e-commerce video creation platform. The availability of user-generated BGM editing logs on this platform allows us to reliably trace back to the source tracks.
In sum, our main contributions are as follows:
@@
  • Evaluation. We adapt and evaluate a diverse set of models, including six VG models (Moment-DETR [16], QD-DETR [23], TR-DETR [26], EaTR [14], and UVCOM [30]), one V2MR model (MVPt [27]), and one Video-corpus Moment Retrieval model (CONQUER [12]). The extensive evaluation highlights the challenging nature of MGSV and establishes MaDe as a stronger baseline.
@@
Figure 2: Visualization of MGSV-EC. Note that the video tag cloud is merely to demonstrate the richness of the visual content. No video tag is used for modeling. Best viewed in color.
@@ 2 Related Work
In developing our solution for MGSV, we draw much inspiration from recent studies on Transformers-based Video Grounding (VG) and Video-to-Music Retrieval (V2MR).
@@ 0 SA-CA×2
VG, aka video moment retrieval by natural language, involves localizing a moment within an untrimmed video that corresponds to a given natural language text. While earlier methods for VG typically rely on pre-generated segments as candidate moments or moment proposals [32, 22, 34], more recent models, equipped with DETR-like decoding modules, directly regress the temporal boundaries in a proposal-free manner [16, 23, 26, 14, 30]. Key elements of these models are summarized in Tab. 1. For instance, consider Moment-DETR [16], the first work that adapted DETR [1] for VG. Given a sequence of visual tokens extracted from video frames and textual tokens from the input text, Moment-DETR first enhances the unimodal features and aligns their dimensions using fully connected (FC) layers. The enhanced tokens are then concatenated along their temporal axis and passed through two self-attention (SA) blocks for multimodal feature fusion. Subsequently, a cross-attention (CA) based decoder is employed, utilizing an array of 10 trainable query tokens as $Q$, with the previously fused tokens serving as $K$ and $V$ for moment location prediction. Note that each query token for a specific CA is obtained as the sum of a positional token, which is shared across all CAs, and a content token from the preceding CA. For its first CA, Moment-DETR initializes the content token with a ZERO vector. As shown in Tab. 1, subsequent works such as QD-DETR [23], TR-DETR [26], and UVCOM [30] primarily focus on innovations in multimodal feature fusion to enhance the interaction between visual and textual tokens, largely through the use of cross-attention (CA) mechanisms. However, in our preliminary experiments, we find that models employing CA-based fusions are more prone to overfitting in the new task. Consequently, we remove CA from the fusion module.
@@
There have been notable efforts to extend VG to a video-corpus scenario, known as Video Corpus Moment Retrieval (VCMR) [12, 33, 31]. In particular, these methods first perform text-to-video retrieval to select the top-$k$ candidate videos from a given corpus. Subsequently, moment detection is conducted for each candidate video. It is worth noting that the above works implement moment detection in a proposal-based manner. For instance, SQuiDNet [31] treats each sampled frame as a candidate point for the start or end timestamp of the moment to be detected, thereby formulating the task as a binary classification problem. Consequently, the temporal resolution of the detection result is limited by the frame sampling frequency.
@@
V2MR aims to retrieve music tracks relevant to a given query video through video-to-music matching. Earlier approaches to V2MR utilized two-branch MLPs for cross-modal feature alignment [11, 24, 2]. In contrast, current methods are based on Transformers, employing self-attention (SA) mechanisms for long-range temporal modeling, see MVPt [27]. Building on MVPt as their backbone, ViML [21] fuses text and video input representations to query music samples, while MuseChat [4] introduces an interactive V2MR system that enhances retrieval results through conversational interactions. Despite these advancements in video-music matching, music grounding is unexplored in these efforts.
@@ 3.1 Dataset Curation
Raw Data Gathering. With permission, we gathered videos posted on an E-commerce video creation platform in the period of Jan. to Sep. 2023, resulting in an initial set of 100k videos. The videos have a frame rate of 34 and a raw resolution of $1080\times 1920$. Each video is associated with a BGM editing log, indicating which music track was used and which part of the music was adopted as the BGM. A set of 9k music tracks were used in total. As the raw video set is quite diverse with varied data quality, automated data cleaning is performed as follows.
Data Cleaning. Given the positive correlation between video content quality and view counts / follows, we empirically exclude videos that have no more than 200 view counts or 100 follows. Videos too short ($<$5 seconds) or too long ($>$50 seconds) are also removed. With the above cleaning procedure, we preserve 53,194 short videos accompanied with 35,393 different music moments from 4,050 unique music tracks. Among the tracks, songs and instrumental music are approximately in a 6:4 ratio. As Fig. 2a shows, the video duration is 23.9 seconds on average, whilst that of the music tracks is 138.9 seconds, indicating a noticeable duration gap between the query videos and the music tracks to be retrieved. Fig. 2e shows a large disparity between the ground-truth moment positions and those detected by a music highlight extractor [^2] https://github.com/remyhuang/pop-music-highlighter/ [13]. These results clearly shows the challenging nature of the new MGSV task.
Data Split. For reproducible research, we recommend data split as follows. A set of 2k videos is randomly sampled to form a held-out test set, while another random set of 2k videos is used as a validation set (for hyper-parameter tuning, model selection, etc.), with the remaining data used for training, see Tab. 2.
@@ 4.1.2 Video-Music Matching
We improve video-text matching by adopting X-Pool [8] to obtain a video-conditioned track embedding $h_{1}(M)$. Originally developed for video-text retrieval, X-Pool aggregates an array of local features with a cross-attention (CA) mechanism. In the current context, the video embedding $h(v)$, obtained by mean pooling, is used as $Q$, whilst $\{\hat{s}_{i}\}_{i=1}^{S}$ are used as $K$ and $V$. The attention weights generated by CA are used for weighted feature aggregation. The weights are computed based on the dot product between $Q$ and $K$ and thus reflect segment-wise relevance w.r.t. the video. Hence, music segments more close to the video will contribute more to $h_{1}(M)$. The final video-music similarity $p_{s}$ is calculated as the sum of the cosine similarity ($cs$) between $h(v)$ and $h_{0}(M)$ and that of $h(v)$ and $h_{1}(M)$. In short, we have X-Pool enhanced video-music matching as:
@@ 5.1 Single-music Grounding
VCMR re-purposed:
@@ 5.2 Music-set Grounding
Human Evaluation. We further conduct a human evaluation on a random subset of 100 test videos to subjectively assess the quality of the MsG results. In particular, each video is individually paired with the music moment predicted by MaDe and its counterpart by MVPt+ / UVCOM+. A subject is asked to watch the two resultant videos (labeled as Video A and Video B without revealing the method used) and, based on their overall visual and audio perception, select one of the following four options: A is better, B is better, neither acceptable, or both acceptable. Our evaluation team comprises seven males and three females, all in their 20s and 30s, with frequent exposure to online short videos and music. Fig. 4 shows that MaDe exceeds the composite solution by approximately 10% (39.4% vs. 28.0%).
Figure 4: Human evaluation results on a random subset of 100 test videos.
5.3 Ablation Study of MaDe
@@
Table 6: MsG performance of MaDe with varied numbers of CAs for decoding.
@@
Table 7: Inference efficiency per mini-batch. Batch size: 512. Raw visual / audio tokens are assumed to be precomputed and thus excluded from efficiency analysis.
@@ 6 Conclusions
Limitations. Our study has clearly shown that CA-based multimodal feature fusion negatively impacts music grounding performance. The surface reason is that such a fusion mechanism tends to cause overfitting. However, the question of how video-music interaction needed for Music Grounding fundamentally differs from the text-video interaction required by Video Grounding remains open.

### v3

@@ Abstract
Adding proper background music helps complete a short video to be shared. Previous work tackles the task by video-to-music retrieval (V2MR), aiming to find the most suitable music track from a collection to match the content of a given query video. In practice, however, music tracks are typically much longer than the query video, necessitating (manual) trimming of the retrieved music to a shorter segment that matches the video duration. In order to bridge the gap between the practical need for music moment localization and V2MR, we propose a new task termed Music Grounding by Short Video (MGSV). To tackle the new task, we introduce a new benchmark, MGSV-EC, which comprises a diverse set of 53k short videos associated with 35k different music moments from 4k unique music tracks. Furthermore, we develop a new baseline method, MaDe, which performs both video-to-music matching and music moment detection within a unified end-to-end deep network. Extensive experiments on MGSV-EC not only highlight the challenging nature of MGSV but also set MaDe as a strong baseline.
Figure 1: Music grounding by short video (MGSV), aiming to localize within a music-track collection a music moment that best serves as background music for the query video. Text in callout boxes to the right of each music moment is for illustrative purposes only.
11footnotetext: Work performed as an intern at Kuaishou.22footnotetext: Corresponding author (xirong@ruc.edu.cn)
@@ 1 Introduction
To address the new task, one might consider applying V2MR at a fine-grained segment level, with music tracks pre-trimmed into shorter segments. This approach is problematic due to the varying durations of videos. While MGSV is new, we observe in a broader context its conceptual resemblance to Video Grounding [16, 14, 30], which aims to localize a moment in a specific video w.r.t. a given textual query. In particular, the query and target modalities in MGSV (Video Grounding) are video (text) and music (video), respectively. Thus, a Video Grounding method could, in principle, be repurposed for MGSV in a single-music mode, assuming that the relevant music track for a query video is provided. When dealing with a collection of music tracks, V2MR is required to identify the relevant track. While combining existing Video Grounding and V2MR methods provides a good starting point, these methods were not originally designed for MGSV. By incorporating task-specific optimizations, we develop a stronger baseline that performs video-to-music Matching and music moment Detection (MaDe) in a single network.
While few public datasets (mainly HIMV [11]) exist for V2MR, each video in these datasets is paired with an equal-length music segment, yet the source tracks from which these segments were derived are absent. Therefore, they cannot be repurposed to MGSV. We construct a new dataset, MGSV-EC, sourced from an e-commerce video creation platform. The availability of user-generated BGM editing logs on this platform allows us to reliably trace back to the source tracks. Noting that music selection for short videos involves inherent subjectivity, our work explores if an automated system can recommend music comparable to that made by skilled practitioners. As such, this work is a starting point for answering the higher-level challenge of diversity. In sum, our main contributions are as follows:
@@
  • Evaluation. We adapt and evaluate a diverse set of models, including six Video Grounding models (Moment-DETR [16], QD-DETR [23], TR-DETR [26], EaTR [14], and UVCOM [30]), one V2MR model (MVPt [27]), and one Video-corpus Moment Retrieval model (CONQUER [12]). The evaluation highlights the challenging nature of MGSV and establishes MaDe as a stronger baseline.
@@
Figure 2: Visualization of MGSV-EC. Note that the video tag cloud is merely to demonstrate the richness of the visual content.
@@ 2 Related Work
In developing our solution for MGSV, we draw inspiration from recent studies on Transformers-based Video Grounding and Video-to-Music Retrieval (V2MR).
@@ 0 SA-CA×2
Video Grounding, aka video moment retrieval by natural language, involves localizing a moment within an untrimmed video that corresponds to a given natural language text. While earlier methods for this task typically rely on pre-generated segments as candidate moments or moment proposals [32, 22, 34], more recent models, equipped with DETR-like decoding modules, directly regress temporal boundaries in a proposal-free manner [16, 23, 26, 14, 30]. Key elements of these models are summarized in Tab. 1. For instance, consider Moment-DETR [16], the first work that adapted DETR [1] for the task. Given a sequence of visual tokens extracted from video frames and textual tokens from the input text, Moment-DETR first enhances the unimodal features and aligns their dimensions using fully connected (FC) layers. The enhanced tokens are then concatenated (ⓒ) along their temporal axis and passed through two self-attention (SA) blocks for multimodal feature fusion. Subsequently, a cross-attention (CA) based decoder is employed, utilizing an array of 10 trainable query tokens as $Q$, with the previously fused tokens serving as $K$ and $V$ for moment location prediction. Note that each query token for a specific CA is obtained as the sum of a positional token, which is shared across all CAs, and a content token from the preceding CA. For its first CA, Moment-DETR initializes the content token with a ZERO vector. As shown in Tab. 1, subsequent works such as QD-DETR [23], TR-DETR [26], and UVCOM [30] primarily focus on innovations in multimodal feature fusion to enhance the interaction between visual and textual tokens, largely through the use of cross-attention (CA) mechanisms. However, in our preliminary experiments, we find that models employing CA-based fusions are more prone to overfitting in the new task. Consequently, we remove CA from the fusion module.
@@
There have been notable efforts to extend video grounding to a video-corpus scenario, known as Video Corpus Moment Retrieval (VCMR) [12, 33, 31]. In particular, these methods first perform text-to-video retrieval to select the top-$k$ candidate videos from a given corpus. Subsequently, moment detection is conducted for each candidate video. It is worth noting that the above works implement moment detection in a proposal-based manner. For instance, SQuiDNet [31] treats each sampled frame as a candidate point for the start or end timestamp of the moment to be detected, thereby formulating the task as a binary classification problem. Consequently, the temporal resolution of the detection result is limited by the frame sampling frequency.
@@
V2MR aims to retrieve music tracks relevant to a given query video through video-to-music matching. Earlier approaches utilized two-branch MLPs for cross-modal feature alignment [11, 24, 2]. In contrast, current methods are based on Transformers, employing self-attention (SA) mechanisms for long-range temporal modeling, see MVPt [27]. Building on MVPt as their backbone, ViML [21] fuses text and video input representations to query music samples, while MuseChat [4] introduces an interactive V2MR system that enhances retrieval results through conversational interactions. Despite these advancements in video-music matching, music grounding is unexplored in these efforts.
@@ 3.1 Dataset Curation
Raw Data Gathering. With permission, we gathered videos posted on an E-commerce video creation platform in the period from Jan. to Sep. 2023, resulting in an initial set of 100k videos. The videos have a frame rate of 34 and a raw resolution of $1080\times 1920$. Each video is associated with a BGM editing log, indicating which music track was used and which part of the music was adopted as the BGM. A set of 9k music tracks was used in total. As the raw video set is quite diverse with varied data quality, automated data cleaning is performed as follows.
Data Cleaning. Given the positive correlation between video content quality and view counts / follows, we empirically exclude videos that have no more than 200 view counts or 100 follows. Videos too short ($<$5 seconds) or too long ($>$50 seconds) are also removed. With the above cleaning procedure, we preserve 53,194 short videos accompanied with 35,393 different music moments from 4,050 unique music tracks. Among the tracks, songs and instrumental music are approximately in a 6:4 ratio. As Fig. 2a shows, the video duration is 23.9 seconds on average, whilst that of the music tracks is 138.9 seconds, indicating a noticeable duration gap between the query videos and the music tracks to be retrieved. Fig. 2e shows a large disparity between the ground-truth moment positions and those detected by a music highlight extractor [^2] https://github.com/remyhuang/pop-music-highlighter/ [13]. These results clearly show the challenging nature of the new MGSV task.
Data Split. For reproducible research, we recommend a data split as follows. A set of 2k videos is randomly sampled to form a held-out test set, while another random set of 2k videos is used as a validation set (for hyper-parameter tuning, model selection, etc.), with the remaining data used for training, see Tab. 2.
@@ 4.1.2 Video-Music Matching
We improve video-text matching by adopting X-Pool [8] to obtain a video-conditioned track embedding $h_{1}(M)$. Originally developed for video-text retrieval, X-Pool aggregates an array of local features with a cross-attention (CA) mechanism. In the current context, the video embedding $h(v)$, obtained by mean pooling, is used as $Q$, whilst $\{\hat{s}_{i}\}_{i=1}^{S}$ are used as $K$ and $V$. The attention weights generated by CA are used for weighted feature aggregation. The weights are computed based on the dot product between $Q$ and $K$ and thus reflect segment-wise relevance w.r.t. the video. Hence, music segments closer to the video will contribute more to $h_{1}(M)$. The final video-music similarity $p_{s}$ is calculated as the sum of the cosine similarity ($cs$) between $h(v)$ and $h_{0}(M)$ and that of $h(v)$ and $h_{1}(M)$. In short, we have X-Pool enhanced video-music matching as:
@@ 5.1 Single-music Grounding
Video Corpus Moment Retrieval re-purposed:
@@ 5.2 Music-set Grounding
Model
SmG
V2MR
MsG
mIoU R1 R5 R10 MoR1 MoR10 MoR100
MVPt+ / UVCOM+ $0.676_{\pm 0.013}$ $7.0_{\pm 0.4}$ $12.5_{\pm 0.7}$ $15.5_{\pm 0.5}$ $5.6_{\pm 0.1}$ $12.0_{\pm 0.3}$ $24.3_{\pm 1.0}$
MaDe $0.743_{\pm 0.017}$ $9.0_{\pm 0.5}$ $17.3_{\pm 0.7}$ $21.1_{\pm 1.0}$ $8.4_{\pm 0.3}$ $18.0_{\pm 0.5}$ $32.7_{\pm 1.4}$
Table 5: Average performance across three distinct data splits.
To verify the robustness of our findings across different data configurations, we conducted extra experiments using two newly generated random splits, following the same data partitioning protocol (Sec. 3.1). The superior performance of MaDe compared to the best baseline remains, see Tab. 5.
Human Evaluation. We further conduct a human evaluation on a random subset of 100 test videos to subjectively assess the quality of the MsG results. In particular, each video is individually paired with the music moment predicted by MaDe and its counterpart by MVPt+ / UVCOM+. A subject is asked to watch the two resultant videos (labeled as Video A and Video B without revealing the method used) and, based on their overall visual and audio perception, select one of the following four options: A is better, B is better, neither acceptable, or both acceptable. Our evaluation team comprised 15 males and 5 females, all in their 20s and 30s, with frequent exposure to online short videos and music. Fig. 4 shows that MaDe exceeds the composite solution by approximately 10% (38.1% vs. 28.6%).
Figure 4: Human evaluation results.
@@ 5.3 Ablation Study of MaDe
Table 7: Numbers of CAs for decoding.
CA for fusion. We conduct a controlled comparison between QD-DETR, UVCOM and their CA-free variants (QD-DETR+, UVCOM+) at varied training data scales. Fig. 5 reveals an inverse relationship: QD-DETR declines with more data, while QD-DETR+ improves. This result suggests that CA-based multimodal feature fusion is the source of incompatibility with the MGSV task’s demands.
@@
Figure 5: SmG performance with varying training data sizes.
@@
Table 8: Inference efficiency per mini-batch. Batch size: 512. Raw visual / audio tokens are precomputed and thus excluded.
@@ 6 Conclusions
Limitations. Our study has clearly shown that CA-based multimodal feature fusion negatively impacts music grounding performance. The surface reason is that such a fusion mechanism tends to cause overfitting. However, the question of how video-music interaction needed for Music Grounding fundamentally differs from the text-video interaction required by Video Grounding remains open. How to customize music grounding with respect to the users’ personal preferences is worthy of future research.
Acknowledgments
This research was supported by NSFC (No.62172420) and Kuaishou. We thank Yingtong Liu, Yuchuan Deng, Yiyi Chen, and Bo Wang for valuable discussion and feedback on this research.